OpenStudy (anonymous):

Prove the Pythagorean Theorem using similar triangles. The Pythagorean Theorem states that in a right triangle, the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse. Be sure to create and name the appropriate geometric figures. How would I start this ???

OpenStudy (anonymous):

OKe

OpenStudy (anonymous):



OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

I messed up on the first image use this one

OpenStudy (anonymous):

Now, by using the Pythagorean theorem, we can plug in two sides, to find a third. Now for the smaller triangle we can find the hypotenuse (5) by plugging in side (4), and side (3).

OpenStudy (anonymous):

\[a^2 + b^ = c^2 \] \[3^2 + 4^2 = c^2\] Now if we work this out we find that \[3^2 + 4^2 = c^2\] \[9 + 16 = c^2\] Now we combine 9 and 16 then get c by itself by squaring everything \[15 = c^2\] \[\sqrt{15} = \sqrt{c^2}\] When you squareroot a square it cancels \[5 = c\]

OpenStudy (anonymous):

Now we just proved that the pythagorean theorem works for the smaller triangle, now all we need to do is enter the corresponding sides for the bigger triangle. Then you're done

OpenStudy (anonymous):

If you need more help mention me with @lugerm, if this is all you needed then please, bestresponse/medal me. Feel free to fan me for future help.

Directrix (directrix):

>Prove the Pythagorean Theorem using similar triangles. This work is not a proof. It is more of a demonstration or argument for one specific pair of similar triangles.

OpenStudy (irishboy123):

https://www.youtube.com/watch?v=pVo6szYE13Y cool music too!!